OpenStudy (learningallday):

Which related equation can be used to solve the equation? Column A Column B 1. –8.7 = k + 6 2. –6 = k + 8.7 3. 8.7 = k + (–6) 4. 6 = k + 8.7 A. k = –8.7 – 6 B. k = 6 – 8.7 C. k = 8.7 – (–6)

OpenStudy (learningallday):

@youngstudier

OpenStudy (learningallday):

i think 1 b 2 c 3a

OpenStudy (youngstudier):

I'm confused.....

OpenStudy (youngstudier):

That's a really weird format....

OpenStudy (learningallday):

Which related equation can be used to solve the equation? Column A Column B 1. –8.7 = k + 6 2. –6 = k + 8.7 3. 8.7 = k + (–6) 4. 6 = k + 8.7 A. k = –8.7 – 6 B. k = 6 – 8.7 C. k = 8.7 – (–6)

OpenStudy (youngstudier):

.....Urm... This is a really screwy question....

OpenStudy (youngstudier):

Can you take a screenshot of the question or no?

OpenStudy (youngstudier):

@BishopPatton

OpenStudy (youngstudier):

@Mahmood786123

OpenStudy (youngstudier):

@texaschic101

OpenStudy (youngstudier):

I'm not sure, sorry

OpenStudy (learningallday):

do you know any one who do

OpenStudy (youngstudier):

Um, @texaschic101

OpenStudy (learningallday):

What value of x makes the equation true? –4.9 = –5.7 + x A. x = –10.6 B. x = –0.8 C. x = 0.8 D. x = 10.6